please help me i dont understand how to do these six problems
1) 1-7x=-20
2) 1+5x=26
3) -6x+7=15
4) -3+4x=-47
5) 3x-2=13
6) -3x-4=-19

Answer by stanbon(75887)

1) 1-7x=-20
Add 7x to both sides
Add 20 to both sides
---
1+20 = 7x
7x = 21
Divide both sides by 7 to get:
x = 3
=============================
2) 1+5x=26
Subtract 1 from both sides to get:
5x = 25
Divide by 5 to get:
x = 5
------------------------------
3) -6x+7=15
Add 6x to both sides
Subtract 15 from both sides
6x = 7-15
6x = -8
Divide both sides by 6
x = -4/3
==============================
4) -3+4x=-47
Add 3 to both sides:
4x = -44
Divide by 4 to get:
x = -11
=============================
5) 3x-2=13
3x = 15
x = 5
===========
6) -3x-4=-19
-3x = -15
x = 5
============
Cheers,
Stan H.
==================
Cheers,
Stan H.